# Germanium target sensed by phonon-mediated kinetic inductance detectors

### v1

@@ Abstract
The BULLKID project developed a cryogenic detector designed to address the mass scaling challenge Cruciani et al. (2022). The detector consists of a monolithic array of 60 cubic silicon absorbers of 0.34 g each carved out of a single 3" diameter 5 mm thick wafer. The phonon sensors are Kinetic Inductance Detectors (KIDs), superconducting resonators featuring intrinsic frequency multiplexing, ease of fabrication and reproducibility Day et al. (2003).
@@
A multi-target particle detector would also provide an advantage in the field of dark matter identification: the ability to test the quadratic scaling of the WIMP-nucleon cross section as a function of the atomic number A is a fundamental tool to confirm a positive evidence of WIMP interactions. The comparison of the signal from germanium and silicon absorbers in the same composite array will also allow to reduce systematic errors Billard and others [APPEC Committee Report] (2022).
In this work we describe the design, fabrication and operation of the first KIDs on a germanium substrate. The performances are in line with those obtained on silicon, making it a promising target material for future experiments with KIDs.
@@
The device has been fabricated using a 550 $\mu$m thick, 2" diameter single side polished germanium wafer at the PTA clean room facility in Grenoble sal . To fabricate the resonators the wafer is initially put under ultra-high vacuum and the native $\text{GeO}_{2}$ on the surface is removed by exposure to argon plasma. Afterwards the polished surface is coated by a 60 nm aluminium film by means of an electron gun evaporator. During the cleaning and coating of the wafer the vacuum is never broken, so that the oxide is not allowed to reform and contaminate the metal-substrate interface.
The removal of the germanium oxide is crucial to achieve a sufficiently high quality of the final device: earlier prototypes, using the standard silicon etching time, were either non-functional or exceedingly noisy. The complete removal of the oxide layer was achieved by increasing the etching time from 3 to 10 minutes, allowing to produce the KIDs with quality comparable to the sensors deposited on silicon substrates.
@@
The pixel design is identical to the sensors of the BULLKID array Cruciani et al. (2022). The active area of each pixel is 4 mm2, made of 105 strips of 1880$\times$20 $\mu$m2 spaced by 8 $\mu$m and arranged to form an inductive meander. The two ends of the meander are terminated to four capacitive fingers whose length is trimmed in order to select the resonant frequency of each pixel (Fig. 1 a); in the current version of the array three pixels are fabricated along a single 2 cm long coplanar waveguide. The sensors were simulated using the SONNET son software: the capacitance has been chosen in order to have the devices resonate at frequencies $1/2\pi\sqrt{LC}$ between 0.7 and 0.8 GHz. The spacing in frequency between the resonators is 25 MHz to minimize electrical cross-talk. In order of increasing resonant frequency, KID-1 and KID-2 are on the same side of the feedline, while KID-3 is on the opposite side (Fig. 1 b).
@@
where $f_{0}$ is the resonant frequency, $1/Q=1/Q_{c}+1/Q_{i}$ is the total quality factor, $Q_{c}$ is the coupling quality factor representing the coupling with the feedline and $Q_{i}$ is the internal quality factor representing internal losses.
@@
The impulsive response of each KID has been characterized as a function of the readout power by firing pulses from the LED lamp at constant energy Cardani et al. (2017). The pulses are processed offline with a matched filter in order to maximize the energy resolution Radeka and Karlovac (1967). The readout power optimizing the signal to noise ratio is -64 dBm at each KID.
@@
The estimation of the responsivity $d\phi/dE$ follows the procedure already employed in Refs. Cardani et al. (2018, 2021); Cruciani et al. (2022); Del Castello (2024). The mean response of the detector to the absorption into the substrate of $N$ photons can be written as $\mu=d\phi/dE\cdot\epsilon N$ where $\epsilon=3.1$ eV is the energy of a single 400 nm photon. The variance of the response is the sum of the variance of the noise $\sigma_{0}^{2}$ and the Poisson’s variance:
@@
$d\phi/{dE}$ [mrad/keV] $\sigma_{0}$ [eV] $\eta$ [%] $\dfrac{A_{\text{KID}}}{A_{\text{TOT}}}$
KID-1 $3.6\pm 0.2$ $380\pm 20$ $2.0\pm 0.2$ 0.12
KID-2 $2.5\pm 0.1$ $450\pm 22$ $2.0\pm 0.2$
KID-3 $2.2\pm 0.2$ $540\pm 31$ $1.6\pm 0.2$
CALDER-17 $5.8$ $115\pm 6$ $7.4-9.4$ 0.42
@@
Table 2: Results of the fit of Eq. 3 to the data acquired by firing LED pulses on the backside of the array. The value of the phonon energy conversion efficiency $\eta$ are derived by inverting Eq. 3. The efficiency is proportional to the ratio of sensitive metal area to the total metal area $A_{\text{KID}}/A_{\text{TOT}}$. Values for CALDER-17 are also reported for comparison.
@@
By inverting Eq. 2 we extract the energy to quasi-particle conversion efficiency $\eta$. The resulting values are summarized in Tab. 2, in the range $(1.6-2.0)\%$. Since the optical fiber is firing on a spot between KID-1 and KID-2 it is expected that they share a similar value while the more distant KID-3 has a reduced efficiency. The energy conversion efficiency is a factor 4 lower than in the CALDER-17 prototype that achieved $(7.4-9.4)\%$ Cardani et al. (2018) depending on the position of the fiber spot. Since the entire metallized structure is able to absorb phonons but only the inductive meander is able to generate a signal, the reduced $\eta$ can be explained as being produced by a competition for phonon collection between KIDs themselves and the inactive metal structures (ie. the feedline and the capacitor) Martinez et al. (2019). The ratio of the active metal surface of a single KID over the total metallized surface was approximately $0.42$ in CALDER-17, while it is only $0.12$ for the sensors discussed in this work. Moreover, for the realization of the current array a single-side polished wafer was used, in place of the double-sided polished wafer for CALDER-17. It is reasonable to believe that an additional source of loss is given by the unpolished surface that could act as a trapping site for phonons.
@@
We can therefore ascribe the difference in baseline resolution with respect to CALDER-17 only to the geometry of the metal layer on the tile, which affects the phonon collection efficiency. There is no evidence that the material used as substrate influences the responsivity and the energy resolution.
@@
In this work we presented the design, fabrication and operation of the first array of KIDs developed on a germanium substrate for phonon-mediated particle detection. The fabrication process was optimized to remove the germanium dioxide from the substrate, as it was found to be a limiting factor for the quality of the resonators. We estimated detection performances which are in line to that obtained with silicon substrates, paving the way to future developments for applications in the neutrino and dark matter fields.
@@ Acknowledgements.
We thank the team of the Laboratorio Rivelatori Criogenici of Sapienza U. for its support. This work was partially supported through the European Research Council through the Consolidator Grant “DANAE” number 101087663. We acknowledge the support of the PTA platform for the fabrication of the device. We thank A. Girardi and M. Iannone of the INFN Sezione di Roma for technical support.

### v2

@@ Abstract
The BULLKID project developed a cryogenic detector designed to address the mass scaling challenge Cruciani et al. (2022). The detector consists of a monolithic array of 60 cubic silicon absorbers of 0.34 g each carved out of a single 3" diameter 5 mm thick wafer. The phonon sensors are Kinetic Inductance Detectors (KIDs), superconducting resonators featuring intrinsic frequency multiplexing, ease and consistency in fabrication Day et al. (2003).
@@
A multi-target particle detector would also provide an advantage in the field of dark matter identification: the ability to test the quadratic scaling of the WIMP-nucleon cross section as a function of the atomic number A is a fundamental tool to confirm a positive evidence of WIMP interactions. The comparison of the signal from germanium and silicon absorbers in the same composite array will also allow to reduce systematic errors Billard et al. (2022).
In this work we describe the design, fabrication and operation of KIDs on a germanium substrate. The performances are in line with those obtained on silicon, making it a promising target material for future experiments with KIDs.
@@
The device has been fabricated using a 550 $\mu$m thick, 2" diameter single side polished germanium wafer at the PTA clean room facility in Grenoble sal . To fabricate the resonators the wafer is initially put under high vacuum and the native $\text{GeO}_{2}$ on the surface is removed by exposure to argon plasma. Afterwards the polished surface is coated by a 60 nm aluminium film by means of an electron gun evaporator. During the cleaning and coating of the wafer the vacuum is never broken, so that the oxide is not allowed to reform and contaminate the metal-substrate interface.
The removal of the germanium oxide is crucial to achieve a sufficiently high quality of the final device: a 10 minutes argon plasma etch achieved a complete removal of the germanium oxide, resulting in KIDs with quality factors and noise comparable to those on silicon substrates. Incomplete removal of the germanium oxide results in greatly reduced internal quality factors (e.g. $Q_{i}<10$ k), as shown by prototypes which used a shorter argon plasma etching time of 3 minutes.
@@
The pixel design is identical to the sensors of the BULLKID array Cruciani et al. (2022). The active area of each pixel is 4 mm2, made of 105 strips of 1880$\times$20 $\mu$m2 spaced by 8 $\mu$m and arranged to form an inductive meander. The two ends of the meander are terminated to four capacitive fingers whose length is adjusted in order to select the resonant frequency of each pixel (Fig. 1 a); in the current version of the array three pixels are fabricated along a single 2 cm long coplanar waveguide. The sensors were simulated using the SONNET son software: the capacitance has been chosen in order to have the devices resonate at frequencies $1/2\pi\sqrt{LC}$ between 0.7 and 0.8 GHz. The spacing in frequency between the resonators is 25 MHz to minimize electrical cross-talk. In order of increasing resonant frequency, KID-1 and KID-2 are on the same side of the feedline, while KID-3 is on the opposite side (Fig. 1 b). The coplanar waveguide geometry has been simulated and optimized for the dielectric constant of germanium $\epsilon_{r}^{Ge}\sim 16$ (while $\epsilon_{r}^{Si}\sim 12$Dunlap and Watters (1953) in order to provide an impedance of $50~{}\Omega$.
@@
where $f_{0}$ is the resonant frequency, $Q=(1/Q_{c}+1/Q_{i})^{-1}$ is the total quality factor, $Q_{c}$ is the coupling quality factor representing the coupling with the feedline and $Q_{i}$ is the internal quality factor representing internal losses.
@@
The impulse response of each KID has been characterized as a function of the readout power by firing pulses from the LED lamp at constant energy Cardani et al. (2017). The pulses are processed offline with a matched filter in order to maximize the energy resolution Radeka and Karlovac (1967). The readout power optimizing the signal to noise ratio is -64 dBm at each KID.
@@
The estimation of the responsivity $d\phi/dE$ follows the procedure validated against a 6 keV X-ray source in Ref. Cardani et al. (2018) and employed in Refs. Cardani et al. (2021); Cruciani et al. (2022); Del Castello (2024). The mean response of the detector to the absorption into the substrate of $N$ photons can be written as $\mu=d\phi/dE\cdot\epsilon N$ where $\epsilon=3.1$ eV is the energy of a single 400 nm photon. The variance of the response is the sum of the variance of the noise $\sigma_{0}^{2}$ and the Poisson’s variance:
@@
[…]
KID-1 1.16 $3.6\pm 0.2$ $380\pm 20$ $2.0\pm 0.2$ 0.12
KID-2 " $2.5\pm 0.1$ $450\pm 22$ $2.0\pm 0.2$
KID-3 " $2.2\pm 0.2$ $540\pm 31$ $1.6\pm 0.2$
CALDER-17 0.28 $5.8$ $115\pm 6$ $7.4-9.4$ 0.42
@@
Table 2: Results of the fit of Eq. 3 to the data acquired by firing LED pulses on the backside of the array. The value of the phonon energy conversion efficiency $\eta$ are derived by inverting Eq. 3. The efficiency is proportional to the ratio of sensitive metal area to the total metal area $A_{\text{KID}}/A_{\text{TOT}}$. Values for CALDER-17 are also reported for comparison, as well as the total mass of the Ge and Si substrates.
By inverting Eq. 2 we extract the energy to quasi-particle conversion efficiency $\eta$. The resulting values are summarized in Tab. 2, in the range $(1.6-2.0)\%$. Since the optical fiber is firing on a spot between KID-1 and KID-2 it is expected that they share a similar value while the more distant KID-3 has a reduced efficiency. The energy conversion efficiency is a factor 4 lower than in the CALDER-17 prototype that achieved $(7.4-9.4)\%$ Cardani et al. (2018) depending on the position of the fiber spot.
@@
The origin of the reduced value of $\eta$ can be understood by comparing the layout of the two devices. The CALDER-17 detector implements a single KID, so that phonons produced can either escape via the substrate supports, be absorbed by the KID or by the insensitive metal (ie. the feedline and the capacitor). In the device of this work, aside the supports and the insensitive metal, the three KIDs share the phonons Martinez et al. (2019). The ratio of the active metal surface of each KID over the total metallized surface (${A_{\text{KID}}}/{A_{\text{TOT}}}$) is approximately $0.42$ in CALDER-17, while it is only $0.12$ for one of the KIDs of this work. Moreover, a single-side polished wafer was used for the realization of the current array, in place of the double-sided polished wafer for CALDER-17. It is reasonable to believe that an additional source of loss is given by the unpolished surface that could act as a trapping site for phonons.
We conclude that there is no evidence that the material used as substrate influences the responsivity and the energy resolution. This seems to be backed up also by the similar values of the expected athermal phonon transmission to Aluminum from Germanium and Silicon. Following the method in Ref. Kaplan et al. (1976), we indeed obtain $\sim 90\%$ and $\sim 80\%$ phonon transmission to Aluminum from Silicon and Germanium, respectively.
@@
We can therefore ascribe the difference in baseline resolution with respect to CALDER-17 only to the geometry of the metal layer on the tile, which affects the phonon collection efficiency. If we assume the value of $\eta=24\%$ of a BULLKID-like geometry, we expect an improvement of the baseline energy resolution of KID-1 down to $380\times 2\%/24\%=32$ eV, in line with the 26 eV of the silicon prototype.
@@
In this work we presented the design, fabrication and operation of the first KIDs developed on a germanium substrate. The fabrication process was optimized to remove the germanium dioxide from the substrate, as it was found to be a limiting factor for the quality of the resonators. We estimated detection performances which are in line to that obtained with silicon substrates, paving the way to future developments for applications in the neutrino and dark matter fields.
@@ Acknowledgements.
We thank the team of the Laboratorio Rivelatori Criogenici of Sapienza U. for its support. This work was partially supported through the European Research Council through the Consolidator Grant “DANAE” number 101087663. We acknowledge the support of the PTA platform for the fabrication of the device. We thank A. Girardi and M. Iannone of the INFN Sezione di Roma for technical support. We thank A. L. De Santis for useful discussions on phonon propagation.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.